# LoRA Done RITE: Robust Invariant Transformation Equilibration for LoRA Optimization

### v1

@@ Abstract
Low-rank adaption (LoRA) is a widely used parameter-efficient finetuning method for LLMs that reduces memory requirements. However, current LoRA optimizers lack transformation invariance, which leads to weight updates that depend on how the two LoRA factors are scaled or rotated. This deficiency leads to inefficient learning and sub-optimal solutions in practice. This paper introduces LoRA-RITE, a novel adaptive matrix preconditioning method for LoRA optimization, which achieves transformation invariance while being computationally efficient. We provide theoretical analysis to demonstrate the benefit of our method and conduct experiments on various LLM tasks with different models including Gemma 2B, 7B, and mT5-XXL. The results demonstrate consistent improvements over existing optimizers. For example, replacing Adam with LoRA-RITE during LoRA fine-tuning of Gemma-2B yields 4.6% accuracy gain on Super-Natural Instructions and 3.5% accuracy gain across other four LLM benchmarks (HellaSwag, ArcChallenge, GSM8K, OpenBookQA).
@@ 1 Introduction
Despite being widely used in practice, we find that applying standard optimizers to LoRA leads to updates that are not “transformation invariant”. By definition of LoRA in (1), the same update ${\bm{Z}}$ can be decomposed in multiple ways, i.e., ${\bm{Z}}={\bm{A}}_{1}{\bm{B}}_{1}^{\top}={\bm{A}}_{2}{\bm{B}}_{2}^{\top}$. Ideally, an optimizer should yield the same update to ${\bm{Z}}$ regardless of the specific factorization. However, commonly used optimizers with diagonal preconditioners like Adam (Kingma & Ba, 2014), Adagrad (Duchi et al., 2011), RMSProp (Tieleman & Hinton, 2012), and even second-order methods like Shampoo (Gupta et al., 2018) and CASPR (Duvvuri et al., 2024), violate this principle when applied to LoRA. This violation not only presents a mathematical inconsistency but also leads to significant inefficiencies during training. In practice, we observe that one LoRA factor often dominates the optimization process, receiving substantial updates while the other remains nearly fixed. Although this can be partially mitigated by some recent proposed approaches such as employing different learning rates for the two factors (Hayou et al., 2024), we ask the question: is there a more principled way to design an optimizer that inherently enforces transformation invariance for LoRA?
@@
  • We propose LoRA-RITE, the first adaptive matrix preconditioning optimizer for LoRA that is transformation-invariant, the property that is lacking for most existing optimizers when applying to LoRA. Theoretically, we provide a convergence analysis for our method.
@@
  • The proposed optimizer leads to significantly improved performance across multiple datasets and architectures. For instance, when applied to the GSM8K (Cobbe et al., 2021) dataset with a Gemma 7B IT model (Gemma Team et al., 2024), LoRA-RITE achieves a 55.50% accuracy rate. This surpasses the widely-used Adam optimizer (Kingma & Ba, 2014) by a substantial margin (48.37%) and even outperforms the second-best optimizer on this dataset, Lamb (You et al., 2020b) (50.64%), by approximately 5%.
@@ 2.1 Definition of Transformation Invariance
As introduced in formulation (1), LoRA adds a low-rank matrix ${\bm{Z}}={\bm{A}}{\bm{B}}^{\top}$ to the original weight matrix ${\bm{W}}$ and learns the LoRA factors ${\bm{A}}\in\mathbb{R}^{m\times r},{\bm{B}}\in\mathbb{R}^{n\times r}$ to minimize the fine-tuning loss. Observe that many different LoRA factors $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$ can represent the same finetuned weight,
@@
When an optimizer is applied to train LoRA, it will produce different updates, $\delta{\bm{A}}_{1},\delta{\bm{B}}_{1}$ or $\delta{\bm{A}}_{2},\delta{\bm{B}}_{2}$, based on the specific parameterization used. Even though $({\bm{A}}_{1},{\bm{B}}_{1})$ and $({\bm{A}}_{2},{\bm{B}}_{2})$ represent the same finetuned weight ${\bm{Z}}$, the updates using different parameterizations can produce different updates to ${\bm{Z}}$. This suggests a serious inconsistency and implies that the update could be suboptimal under some parameterizations.
Based on this observation, we propose that LoRA optimization should ensure transformation invariance, defined as follows:
@@
Let $({\bm{A}}_{1},{\bm{B}}_{1})$ and $({\bm{A}}_{2},{\bm{B}}_{2})$ be any two pairs of LoRA factors that satisfy (2). An optimizer exhibits transformation invariance if its updates, $(\delta{\bm{A}}_{1},\delta{\bm{B}}_{1})$ and $(\delta{\bm{A}}_{2},\delta{\bm{B}}_{2})$, satisfy
@@
This means the optimizer should produce the same update, $\delta{\bm{Z}}$, to the fine-tuned weights for any equivalent LoRA factorizations. To satisfy the condition (3), the following equality should hold
$\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}=\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}+{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}.$ (4)
The following condition, which will be used in our later derivations, is sufficient for (4) to hold:
$\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}=\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top},\ {\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}={\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top},\ \delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}=\delta{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}.$ (5)
As a special case, scalar scale invariance as introduced in Definition 2 is a weaker version of transformation invariance, which only requires that updates remain equivalent when the LoRA factors are scaled up or down by a constant factor. Formally, we define it as:
@@
Let $({\bm{A}}_{1},{\bm{B}}_{1})$ be a pair of LoRA factors and let ${\bm{A}}_{2}=s{\bm{A}}_{1},{\bm{B}}_{2}=(1/s){\bm{B}}_{1}$ for some nonzero constant $s$. An optimizer exhibits scalar scale invariance if its updates, $(\delta{\bm{A}}_{1},\delta{\bm{B}}_{1})$ and $(\delta{\bm{A}}_{2},\delta{\bm{B}}_{2})$, satisfy
@@
Surprisingly, we will show that most commonly used optimizers, when applied to LoRA, do not even satisfy this weaker form of transformation invariance.
@@ 2.2 Existing Optimizers are not Scalar Scale Invariant
We now show that neither gradient descent nor Adam is scalar scale invariant (2), and in fact, almost all the existing optimizers are not scalar scale invariant when applied to LoRA optimization.
Let us consider the following example:
${\bm{A}}_{2}=s{\bm{A}}_{1},\quad{\bm{B}}_{2}=(1/s){\bm{B}}_{1},$ (6)
then by using chain rule on ${\bm{Z}}={\bm{A}}_{1}{\bm{B}}_{1}^{\top}={\bm{A}}_{2}{\bm{B}}_{2}^{\top}$, we have
${\nabla}{\bm{A}}_{1}=\nabla{\bm{Z}}{\bm{B}}_{1},\text{ and }{\nabla}{\bm{A}}_{2}$ $=\nabla{\bm{Z}}{\bm{B}}_{2}.$ (7)
We can rewrite $\nabla{\bm{A}}_{2}$ as:
$\nabla{\bm{A}}_{2}$ $=(1/s)\nabla{\bm{Z}}{\bm{B}}_{1}\quad(\text{since }{\bm{B}}_{2}=(1/s){\bm{B}}_{1})$
$=(1/s)\nabla{\bm{A}}_{1}.$ (8)
To analyze gradient descent, we set the following in condition (4):
$\delta{\bm{A}}_{1}\coloneqq-\nabla{\bm{A}}_{1},\ \delta{\bm{A}}_{2}\coloneqq-\nabla{\bm{A}}_{2},$ $\ \delta{\bm{B}}_{1}\coloneqq-\nabla{\bm{B}}_{1}\text{ and }\delta{\bm{B}}_{2}\coloneqq-\nabla{\bm{B}}_{2}$
The first term on right hand side of condition (4) can be rewritten as follows:
$\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}$ $=-{\nabla}{\bm{A}}_{2}{\bm{B}}_{2}^{\top}$
[…]
$=-(1/s^{2}){\nabla}{\bm{A}}_{1}{\bm{B}}_{1}^{\top},\quad(\text{since }{\bm{B}}_{2}=(1/s){\bm{B}}_{1})$
$=(1/s^{2})\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}.$
Thus the first terms on both sides of condition (4) have a scale difference of $1/s^{2}$. Similarly, we can show that the second terms have a scale difference of $s^{2}$, while the third terms remain identical. Therefore, gradient descent is not scalar scale invariant, and the gradient can be arbitrary large when $s$ goes to 00 or infinity.
@@
$\delta{\bm{A}}_{1}:=-\nabla{\bm{A}}_{1}/(\nabla{\bm{A}}_{1}\odot\nabla{\bm{A}}_{1})^{1/2}$
be the Adam update for ${\bm{A}}_{1}$, and define updates of ${\bm{A}}_{2},\ {\bm{B}}_{1}$ and ${\bm{B}}_{2}$ similarly. It is worth noting that all the operations above are element-wise and we omit the momentum term for brevity. We have
$\delta{\bm{A}}_{2}=-\frac{\nabla{\bm{A}}_{2}}{(\nabla{\bm{A}}_{2}\odot\nabla{\bm{A}}_{2})^{1/2}}=-\frac{(1/s)\nabla{\bm{A}}_{1}}{(1/s)(\nabla{\bm{A}}_{1}\odot\nabla{\bm{A}}_{1})^{1/2}}=\delta{\bm{A}}_{1}.$
As a result,
$\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}=\delta{\bm{A}}_{1}{\bm{B}}_{2}^{\top}=(1/s){\delta}{\bm{A}}_{1}{\bm{B}}_{1}^{\top}.$
Similar to the analysis for gradient descent, we find a discrepancy in scale of the terms in condition (4) for Adam, thus failing to satisfy scalar scale invariance. Actually, one can see that most of the existing optimizers, such as Adagrad (Duchi et al., 2011), RMSProp (Tieleman & Hinton, 2012), and Shampoo (Gupta et al., 2018) are not scale or transformation invariant.
@@ 2.3 Benefits of Transformation Invariance
Why is transformation invariance important? Beyond the mathematical argument that different parameterizations of the same weight update should be equivalent, we demonstrate that transformation invariance leads to more efficient feature learning.
The concept of efficient feature learning, introduced in (Hayou et al., 2024), describes the asymptotic training behavior of LoRA as the network width grows. As discussed earlier, for LoRA, the update to the matrix ${\bm{Z}}={\bm{A}}{\bm{B}}^{\top}$ can be decomposed into three parts
$\delta{\bm{Z}}$ $=({\bm{A}}+\delta{\bm{A}})({\bm{B}}^{\top}+\delta{\bm{B}}^{\top})-{\bm{A}}{\bm{B}}^{\top}$
$=\delta{\bm{A}}{\bm{B}}^{\top}+{\bm{A}}\delta{\bm{B}}^{\top}+\delta{\bm{A}}\delta{\bm{B}}^{\top},$
where the third term is typically negligible as it depends on the square of the learning rate. Efficient feature learning requires that both $\delta{\bm{A}}{\bm{B}}^{\top}{\bm{x}}$ and ${\bm{A}}\delta{\bm{B}}^{\top}{\bm{x}}$ are of magnitude $\mathcal{O}(n^{0})=\mathcal{O}(1)$ with respect to the network width $n$, where ${\bm{x}}$ is the input embedding. In other words, let the scale be $\mathcal{O}(n^{\alpha})$, then it neither explodes ($\alpha>0$) nor diminishes ($\alpha<0$), when the network width $n$ grows.
Hayou et al. (2024) show that conventional optimizers do not satisfy efficient feature learning. This can be seen from Figure 1, where the weight norm for factor ${\bm{B}}$ changes significantly while the weight norm for factor ${\bm{A}}$ barely changes.
Under mild assumptions, we can show that a transformation-invariant optimizer guarantees efficient feature learning. The proof is given in Appendix.
@@
Beyond the efficient learning guarantee, in practice when training LoRA with existing optimizers, it is often the case that only one of the LoRA factors is updating properly, while the other remains almost unchanged, as shown in Figure 1. This is also a consequence of lacking scalar scale invariance, as the initial scale for the two LoRA factors can be very different (one initialized from 00 while other from Gaussian random).
(a) Weight norm of the ${\bm{A}}$ factor
(b) Weight norm of the ${\bm{B}}$ factor
Figure 1: The weight norm of LoRA factors ${\bm{A}}$ and ${\bm{B}}$ across different training steps for existing baselines, Adam, ScaledAdam and Shampoo, where only one LoRA factor visibly changes.

### v2

@@ Abstract
Low-rank adaption (LoRA) is a widely used parameter-efficient finetuning method for LLMs that reduces memory requirements. However, current LoRA optimizers lack transformation invariance, which leads to weight updates that depend on how the two LoRA factors are scaled or rotated. This deficiency leads to inefficient learning and sub-optimal solutions in practice. This paper introduces LoRA-RITE, a novel adaptive matrix preconditioning method for LoRA optimization, which achieves transformation invariance while being computationally efficient. We provide theoretical analysis to demonstrate the benefit of our method and conduct experiments on various LLM tasks with different models including Gemma 2B, 7B, and mT5-XXL. The results demonstrate consistent improvements over existing optimizers. For example, replacing Adam with LoRA-RITE during LoRA fine-tuning of Gemma-2B yields 4.6% accuracy gain on Super-Natural Instructions and 3.5% accuracy gain across four other LLM benchmarks (HellaSwag, ArcChallenge, GSM8K, OpenBookQA).
@@ 1 Introduction
Despite being widely used in practice, we find that applying standard optimizers to LoRA leads to updates that are not “transformation invariant”. By definition of LoRA in (1), the same update ${\bm{Z}}$ can be decomposed in multiple ways, i.e., ${\bm{Z}}={\bm{A}}_{1}{\bm{B}}_{1}^{\top}={\bm{A}}_{2}{\bm{B}}_{2}^{\top}$. Ideally, an optimizer should yield the same update to ${\bm{Z}}$ regardless of the specific factorization. However, commonly used optimizers with diagonal preconditioners like Adam (Kingma & Ba, 2014), Adagrad (Duchi et al., 2011), RMSProp (Tieleman & Hinton, 2012), and even second-order methods like Shampoo (Gupta et al., 2018) and CASPR (Duvvuri et al., 2024), violate this principle when applied to LoRA. This violation not only presents a mathematical inconsistency but also leads to significant inefficiencies during training. In practice, we observe that one LoRA factor often dominates the optimization process, receiving substantial updates while the other remains nearly fixed. Although this can be partially mitigated by some recently proposed approaches such as employing different learning rates for the two factors (Hayou et al., 2024), we ask the question: is there a more principled way to design an optimizer that inherently enforces transformation invariance for LoRA?
@@
  • We propose LoRA-RITE, the first adaptive matrix preconditioning optimizer for LoRA that is transformation-invariant, the property that is lacking in most existing optimizers when applied to LoRA. Theoretically, we provide a convergence analysis for our method.
@@
  • The proposed optimizer leads to significantly improved performance across multiple datasets and architectures. For instance, when applied to the GSM8K (Cobbe et al., 2021) dataset with a Gemma 7B IT model (Gemma Team et al., 2024), LoRA-RITE achieves a 55.50% accuracy rate. This surpasses the widely-used Adam optimizer (Kingma & Ba, 2014) by a substantial margin (48.37%) and even outperforms the second-best optimizer on this dataset, Lamb (You et al., 2020) (50.64%), by approximately 5%.
@@ 2.1 Definition of Transformation Invariance
As introduced in (1), LoRA adds a low-rank matrix ${\bm{Z}}={\bm{A}}{\bm{B}}^{\top}$ to the original weight matrix ${\bm{W}}$ and learns the LoRA factors ${\bm{A}}\in\mathbb{R}^{m\times r},{\bm{B}}\in\mathbb{R}^{n\times r}$ to minimize the fine-tuning loss. Observe that many different LoRA factors $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$ can represent the same finetuned weight,
@@
When an optimizer is applied to train LoRA, it will produce different updates, $\delta{\bm{A}}_{1},\delta{\bm{B}}_{1}$ or $\delta{\bm{A}}_{2},\delta{\bm{B}}_{2}$, based on the specific parameterization used. Even though $({\bm{A}}_{1},{\bm{B}}_{1})$ and $({\bm{A}}_{2},{\bm{B}}_{2})$ represent the same finetuned weight ${\bm{Z}}$, the updates using different parameterizations can produce different updates to ${\bm{Z}}$. This suggests a serious inconsistency and implies that the update could be suboptimal under some parameterizations. Based on this observation, we propose that LoRA optimization should ensure transformation invariance, defined as follows:
@@
Let $({\bm{A}}_{1},{\bm{B}}_{1})$ be a pair of LoRA factors and let ${\bm{A}}_{2}={\bm{A}}_{1}{\bm{R}},{\bm{B}}_{2}={\bm{B}}_{1}{\bm{R}}^{-\top}$ for some invertible matrix ${\bm{R}}$. An optimizer exhibits transformation invariance if its updates, $(\delta{\bm{A}}_{1},\delta{\bm{B}}_{1})$ and $(\delta{\bm{A}}_{2},\delta{\bm{B}}_{2})$, satisfy
@@
This means the optimizer should produce the same update, $\delta{\bm{Z}}$, to the fine-tuned weights for any equivalent LoRA factorizations. As a special case, we introduce scalar scale invariance below in Definition 2 as a weaker version of transformation invariance, which only requires that updates remain equivalent when the LoRA factors are scaled up or down by a scalar factor. Formally, we define it as:
@@
Let $({\bm{A}}_{1},{\bm{B}}_{1})$ be a pair of LoRA factors and let ${\bm{A}}_{2}=s{\bm{A}}_{1},{\bm{B}}_{2}=(1/s){\bm{B}}_{1}$ for some nonzero scalar constant $s$. An optimizer exhibits scalar scale invariance if its updates, $(\delta{\bm{A}}_{1},\delta{\bm{B}}_{1})$ and $(\delta{\bm{A}}_{2},\delta{\bm{B}}_{2})$, satisfy
@@
Surprisingly, we will show that most commonly used LoRA optimizers do not even satisfy this weaker form of transformation invariance.
@@ 2.2 Existing Optimizers are not Scalar Scale Invariant
We now show that neither gradient descent nor Adam are scalar scale invariant, and in fact, almost all the existing optimizers are not scalar scale invariant when applied to LoRA optimization.
Assume loss is $f({\bm{Z}})$, and the gradient of the loss function $f$ with respect to ${\bm{Z}}$, ${\bm{A}}_{1}$, ${\bm{B}}_{1}$, ${\bm{A}}_{2}$, ${\bm{B}}_{2}$ to be ${\nabla}{\bm{Z}}:={\partial f}/{\partial{\bm{Z}}}$, ${\nabla}{\bm{A}}_{1}:={\partial f}/{\partial{\bm{A}}_{1}}$, ${\nabla}{\bm{B}}_{1}:={\partial f}/{\partial{\bm{B}}_{1}}$, ${\nabla}{\bm{A}}_{2}:={\partial f}/{\partial{\bm{A}}_{2}}$, and ${\nabla}{\bm{B}}_{2}:={\partial f}/{\partial{\bm{B}}_{2}}$ respectively. Recall ${\bm{Z}}={\bm{A}}_{1}{\bm{B}}_{1}^{\top}={\bm{A}}_{2}{\bm{B}}_{2}^{\top}$ and by chain rule, we have
${\nabla}{\bm{A}}_{1}:={\partial f}/{\partial{\bm{A}}_{1}}={\partial f}/{\partial{\bm{Z}}}*{\partial{\bm{Z}}}/{\partial{\bm{A}}_{1}}={\partial f}/{\partial{\bm{Z}}}*{\bm{B}}_{1}=\nabla{\bm{Z}}{\bm{B}}_{1},$ (4)
Similarly, we have ${\nabla}{\bm{B}}_{1}=\nabla{\bm{Z}}^{\top}{\bm{A}}_{1}$,${\nabla}{\bm{A}}_{2}=\nabla{\bm{Z}}{\bm{B}}_{2}$, and ${\nabla}{\bm{B}}_{2}=\nabla{\bm{Z}}^{\top}{\bm{A}}_{2}$. For gradient descent,
$\delta{\bm{A}}_{1}\coloneqq-\eta\nabla{\bm{A}}_{1},~{}\delta{\bm{B}}_{1}\coloneqq-\eta\nabla{\bm{B}}_{1},~{}\delta{\bm{A}}_{2}\coloneqq-\eta\nabla{\bm{A}}_{2},~{}\delta{\bm{B}}_{2}\coloneqq-\eta\nabla{\bm{B}}_{2},$
where $\eta$ is the learning rate. To test scalar scale invariant, let ${\bm{A}}_{2}=s{\bm{A}}_{1},{\bm{B}}_{2}=(1/s){\bm{B}}_{1}$, we have
$\delta{\bm{A}}_{2}$ $=-\eta\nabla{\bm{Z}}{\bm{B}}_{2}=-(1/s)\eta\nabla{\bm{Z}}{\bm{B}}_{1}=(1/s)\delta{\bm{A}}_{1}\quad(\text{since }{\bm{B}}_{2}=(1/s){\bm{B}}_{1})$ (5)
$\delta{\bm{B}}_{2}$ $=-\eta\nabla{\bm{Z}}^{\top}{\bm{A}}_{2}=-s\eta\nabla{\bm{Z}}^{\top}{\bm{A}}_{1}=s\delta{\bm{B}}_{1}\quad(\text{since }{\bm{A}}_{2}=s{\bm{A}}_{1}).$ (6)
Consequently, from (5) and (6):
${\bm{A}}_{2}{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}+{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}$ $={\bm{A}}_{1}{\bm{B}}_{1}^{\top}+(1/s^{2})\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+s^{2}{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}$
$\neq{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top},$
and so (3) does not hold for arbitrary $s$. Therefore, gradient descent is not scalar scale invariant, and the gradient can be arbitrary large for the LoRA factors when $s$ goes to 00 or infinity.
@@
$\delta{\bm{A}}_{1}:=-\eta\nabla{\bm{A}}_{1}/(\nabla{\bm{A}}_{1}\odot\nabla{\bm{A}}_{1})^{1/2}$
be the Adam update for ${\bm{A}}_{1}$, where $\odot$ denotes elementwise multiplication. Note that we omit the momentum term for brevity. Defining updates of ${\bm{A}}_{2},\ {\bm{B}}_{1}$ and ${\bm{B}}_{2}$ similarly, we have
$\delta{\bm{A}}_{2}=-\frac{\eta\nabla{\bm{A}}_{2}}{(\nabla{\bm{A}}_{2}\odot\nabla{\bm{A}}_{2})^{1/2}}=-\frac{(1/s)\eta\nabla{\bm{A}}_{1}}{(1/s)(\nabla{\bm{A}}_{1}\odot\nabla{\bm{A}}_{1})^{1/2}}=\delta{\bm{A}}_{1}$
and similarly $\delta{\bm{B}}_{2}=\delta{\bm{B}}_{1}$. As a result,
${\bm{A}}_{2}{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}+{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}$ $={\bm{A}}_{1}{\bm{B}}_{1}^{\top}+(1/s)\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+s{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}$
$\neq{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top},$
thus failing to satisfy scalar scale invariance. Actually, one can see that most of the existing optimizers, such as Adagrad (Duchi et al., 2011), RMSProp (Tieleman & Hinton, 2012), and Shampoo (Gupta et al., 2018) are not scalar scale or transformation invariant.
@@ 2.3 Benefits of Transformation Invariance
Why is transformation invariance important? Beyond the mathematical argument that different parameterizations of the same weight update should be equivalent, we demonstrate that transformation invariance leads to more efficient feature learning. The concept of efficient feature learning, introduced in (Hayou et al., 2024), describes the asymptotic training behavior of LoRA as the network width grows. As discussed earlier, for LoRA, the update to the matrix ${\bm{Z}}={\bm{A}}{\bm{B}}^{\top}$ can be decomposed into three parts
$\delta{\bm{Z}}$ $=({\bm{A}}+\delta{\bm{A}})({\bm{B}}^{\top}+\delta{\bm{B}}^{\top})-{\bm{A}}{\bm{B}}^{\top}=\delta{\bm{A}}{\bm{B}}^{\top}+{\bm{A}}\delta{\bm{B}}^{\top}+\delta{\bm{A}}\delta{\bm{B}}^{\top},$
where the $\delta{\bm{A}}\delta{\bm{B}}^{\top}$ is typically negligible as it depends on the square of the learning rate. Efficient feature learning requires that both $\delta{\bm{A}}{\bm{B}}^{\top}{\bm{x}}$ and ${\bm{A}}\delta{\bm{B}}^{\top}{\bm{x}}$ are of magnitude $\mathcal{O}(n^{0})=\mathcal{O}(1)$ with respect to the network width $n$, where ${\bm{x}}$ is the input embedding. On the contrary, if the magnitude of $\delta{\bm{A}}{\bm{B}}^{\top}{\bm{x}}$ is dependent on $n$: $\mathcal{O}(n^{\alpha})$, then it can explode if $\alpha>0$ and diminish if $\alpha<0$, as the network width $n$ grows.
Hayou et al. (2024) show that conventional optimizers do not satisfy efficient feature learning. This can be seen from Figure 1, where the weight norm for factor ${\bm{B}}$ changes significantly while the weight norm for factor ${\bm{A}}$ barely changes. More discussions about the this unbalanced training dynamic are in Appendix A.8.
Under mild assumptions, we can show that a transformation-invariant optimizer guarantees efficient feature learning. The proof is given in Appendix A.3.
@@
Beyond the efficient learning guarantee, in practice when training LoRA with existing optimizers, it is often the case that only one of the LoRA factors is updated properly, while the other remains almost unchanged, as shown in Figure 1. This is also a consequence of lacking scalar scale invariance, as the initial scales for the two LoRA factors can be very different (one it typically initialized from 00 while other from Gaussian random).
(a) Frobenius norm of LoRA factor ${\bm{A}}$
(b) Frobenius norm of LoRA factor ${\bm{B}}$
Figure 1: The norms of the LoRA factors ${\bm{A}}$ and ${\bm{B}}$ change differently across different training steps.